OpenStudy (sleepyjess):

Use mathematical induction to prove the statement is true for all positive integers n. The integer n^3 + 2n is divisible by 3 for every positive integer n.

mathslover (mathslover):

First put n = 1 and find whether it is true or not. Then put n = k and assume it is true. Find f(k+1) and prove the required thing.

OpenStudy (sleepyjess):

I did n = 1, but do I have to go further than that or just n = 1?

mathslover (mathslover):

You've to move on to step 2nd.

mathslover (mathslover):

Say, \(f(n) = n^3 + 2n\) i) Step - 1 : \(f(1) = 1^3 + 2(1) = ? \) Find whether the value "?" is divisible by 3 or not. ii) Step - 2 : Assume that the statement holds true for n = k. That is, f(k) is true. That is, \(f(k) = k^3 + 2k = 3m \) where m is any integer.

mathslover (mathslover):

After this, you will find f(k+1) : \(f(k+1) = (k+1)^3 + 2(k+1) \) Use identities for \((a+b)^3\) and \(k^3 + 2k = 3m\) (we assumed this in Step 2nd) Try to find f(k+1) as multiple of 3. You will be done then.

OpenStudy (sleepyjess):

oh! that seems simple enough :)

mathslover (mathslover):

:) Good Luck @sleepyjess !

OpenStudy (sleepyjess):

Thank you so much for helping :)